# EVDA: Evolving Deepfake Audio Detection Continual Learning Benchmark

### v1

@@ Abstract
The rise of advanced large language models such as GPT-4, GPT-4o, and the Claude family has made fake audio detection increasingly challenging. Traditional fine-tuning methods struggle to keep pace with the evolving landscape of synthetic speech, necessitating continual learning approaches that can adapt to new audio while retaining the ability to detect older types. Continual learning, which acts as an effective tool for detecting newly emerged deepfake audio while maintaining performance on older types, lacks a well-constructed and user-friendly evaluation framework. To address this gap, we introduce EVDA, a benchmark for evaluating continual learning methods in deepfake audio detection. EVDA includes classic datasets from the Anti-Spoofing Voice series, Chinese fake audio detection series, and newly generated deepfake audio from models like GPT-4 and GPT-4o. It supports various continual learning techniques, such as Elastic Weight Consolidation (EWC), Learning without Forgetting (LwF), and recent methods like Regularized Adaptive Weight Modification (RAWM) and Radian Weight Modification (RWM). Additionally, EVDA facilitates the development of robust algorithms by providing an open interface for integrating new continual learning methods.
@@ 2.1 Data Architecture
Our benchmark architecture consists of twelve experiences (Exp_1 to Exp_12), each representing different datasets and conditions for detecting synthetic audio. These experiences vary in data sources, language, acoustic conditions, and the number of training and evaluation samples. The data architecture of EVDA benchmark are listed as Table 1.
Table 1: The Data architecture of the Evda deepfake audio detection benchmark. T: True; F: False
Task $\mathbf{Task_{1}}$ $\mathbf{Task_{2}}$ $\mathbf{Task_{3}}$ $\mathbf{Task_{4}}$
Source ASVspoof2015 ASVspoof2019 FoR HAD
Sub Source - LA - -
Have Vocoder ID T T F F
Language English English English Chinese
Acoustic Condition No significant noise Logical access Human Human/Partially fake
Training Samples 2000 2000 2000 2000
Testing Samples 500 500 500 500
Task $\mathbf{Task_{5}}$ $\mathbf{Task_{6}}$ $\mathbf{Task_{7}}$ $\mathbf{Task_{8}}$
Source ASVspoof2021 WaveFake FMFCC-A ADD 2022
Sub Source LA - - -
Have Vocoder ID T T F F
Language English English/Japanese Chinese Chinese
Acoustic Condition Logical access Human Human Low-quality/Partially fake
Training Samples 2000 2000 2000 2000
Testing Samples 500 500 500 500
Task $\mathbf{Task_{9}}$ $\mathbf{Task_{1}0}$ $\mathbf{Task_{11}}$ $\mathbf{Task_{12}}$
Source In-the-Wild ADD 2023 GPT-4 GPT-4o
Sub Source - - - -
Have Vocoder ID F F F F
Language English Chinese English English
Acoustic Condition Real world Low-quality/Partially fake Logical access Logical access
Training Samples 2000 2000 2000 2000
Testing Samples 500 500 500 500
The proposed benchmark architecture, as outlined in the table, is designed to comprehensively evaluate the performance of audio spoofing and anti-spoofing models across a diverse set of conditions and datasets. Each experiment (Exp_1 to Exp_12) has been carefully crafted to test various aspects of the detection models under different acoustic and linguistic environments, source integrity, and data authenticity. Below, we detail each experiment’s setup and objectives:
  • Exp_1: ASVspoof2015 (English)(Wu et al., 2015) - Utilizes 2000 training and 500 evaluation samples from the ASVspoof2015 dataset, focusing on English language audio in clean conditions without significant channel or background noise effects. Vocoder ID is available, aiding in more precise synthetic speech detection.
  • Exp_2: ASVspoof2019 - Logical Access (LA) (English)(Todisco et al., 2019) - Also incorporates 2000 training and 500 evaluation samples but focuses on logical access scenarios to test models against sophisticated spoofing attacks in English. Vocoder ID is again available, providing critical data for assessing synthetic signal traits.
  • Exp_3 and Exp_4: FoR (Reimao & Tzerpos, 2019) and HAD (Yi et al., 2021) (English and Chinese) - Each uses 2000 training and 500 evaluation samples. FoR involves purely human-generated audio in English, while HAD includes a mix of human and partially fake audio in Chinese, presenting challenges in cross-lingual and authenticity discernment.
  • Exp_5: ASVspoof2021 - Logical Access (English)(Liu et al., 2023) - Continues the logical access examination with the latest dataset iteration, using English audio samples. The presence of vocoder ID helps in identifying new spoofing techniques that have emerged since the 2019 dataset.
  • Exp_6: WaveFake (English/Japanese)(Frank & Schönherr, 2021) - Addresses bilingual detection scenarios, testing the robustness of models against spoofs in both English and Japanese. This dataset’s inclusion highlights the benchmark’s focus on multilingual and cross-cultural spoofing detection.
@@
  • Exp_7 and Exp_8: FMFCC-A (Zhang et al., 2021) and ADD 2022 (Yi et al., 2022) (Chinese) - Each experiment uses 2000 training and 500 evaluation samples of Chinese audio. These tests are critical for models to adapt to low-quality, partially fake audio environments and a generation & detection game setting, increasing the difficulty of the spoofing detection task.
@@
  • Exp_9 and Exp_10: In-the-Wild (Müller et al., 2022) and ADD 2023 (Yi et al., 2023) (English and Chinese) - These experiments push the boundaries by including real-world audio scenarios (Exp_9) and the latest generation & detection games in a low-quality, partially fake setting (Exp_10), testing the models’ effectiveness in dynamically changing and challenging acoustic environments.
  • Exp_11: Collected Audio Generated by GPT-4 (English) (OpenAI, 2023) - This experiment involves 2000 training and 500 evaluation samples generated by the GPT-4 model, tailored to produce synthetic speech that mimics human conversational patterns in English. The experiment aims to evaluate the effectiveness of detection models against the latest AI-generated audio, challenging them to distinguish between human-like synthetic speech and authentic human speech. The sophisticated natural language generation capabilities of GPT-4 represent a significant leap in the realism of generated audio, thus presenting a unique challenge for anti-spoofing technologies.
  • Exp_12: Collected Audio Generated by GPT-4o (English) - Similar to Exp_11, this experiment also uses 2000 training and 500 evaluation samples, but the audio is generated by GPT-4o, an optimized version of GPT-4 designed to handle English languages with enhanced contextual understanding and expression. This experiment tests the models’ abilities to detect synthetic speech generated by GPT-4o, thereby assessing their adaptability and robustness in a globally interconnected and linguistically diverse world.
@@
The benchmark is designed not only to assess the current state-of-the-art in spoofing detection but also to push the development of more robust, adaptable, and cross-lingual models capable of handling the complexities of real-world audio spoofing scenarios.
@@ 2.2 Data Selection Methodology
For each task within our benchmark, we employ two strategies for data selection:
  1. 1.
    Random Selection: We randomly select a fixed number of authentic and spoofed audio samples from the original datasets or collected data. Initially, we split the data into two groups: Real and Fake. From the ”Real” group, we randomly select 1000 real samples, and from the ”Fake” group, we select 1000 fake samples corresponding to various generation methods. These samples are combined to form the training set. For the evaluation set, we follow the same procedure but select only 250 real and 250 fake samples (1/4 of the training set size). This approach ensures a balanced representation of real and fake samples in both the training and evaluation sets.
  2. 2.
@@
    Here, $N_{\text{real}}$ and $N_{\text{fake}}$ represent the number of models classifying the sample as real or fake, respectively. According to information theory (Shannon, 1948), samples with the highest entropy values contain the most information, indicating uncertainty and hence, difficulty in classification. We list the entropy of each sample and rank them from highest to lowest. The top 250 real and 250 fake samples are selected to construct the evaluation set, ensuring it includes the most challenging and informative samples. The next 1000 real and 1000 fake samples are selected for the training set, which benefits from the inclusion of diverse and informative samples.
2.3 Continuous Learning Methods in Benchmark
The benchmark incorporates a variety of continuous learning approaches to evaluate and enhance the adaptability of audio spoofing detection systems. These methods include:
  • CWRStar(Maltoni & Lomonaco, 2019), Replay(Rolnick et al., 2019), Joint Training(Parisi et al., 2019): Basic strategies for integrating new knowledge while retaining previously learned information without severe forgetting.
  • GDumb (Finetuning)(Prabhu et al., 2020), (Bias)Reservoir(Vitter, 1985), Cumulative: Techniques focusing on selectively updating the model or maintaining a subset of previous data to balance between old and new learning.
  • LwF(Li & Hoiem, 2017), OWM(Zeng et al., 2019), RAWM(Zhang et al., 2023), RWM(Zhang et al., 2024), GEM(Lopez-Paz & Ranzato, 2017), AGEM(Chaudhry et al., 2019): Methods that leverage gradients or error restrictions to mitigate catastrophic forgetting.
  • EWC(Kirkpatrick et al., 2017), Finetune, ICaRL(Rebuffi et al., 2017), SIZenke et al. (2017), AR1, StreamingLDA: Advanced strategies that introduce regularization, importance sampling, or dynamically adjust model architecture based on incoming data streams.
In summary, our benchmark is designed to rigorously assess the effectiveness of various continuous learning methods in the context of audio spoofing detection. By incorporating a diverse set of strategies, ranging from basic integration techniques to advanced regularization and dynamic adjustment methods, we aim to provide a comprehensive evaluation framework. This approach not only facilitates the identification of strengths and weaknesses of each method but also paves the way for future advancements in the continual learning domain. The ultimate goal is to develop robust and adaptive models that can efficiently handle the evolving nature of audio spoofing attacks, ensuring high reliability and accuracy in real-world applications.
@@ 3 Result
We evaluate our benchmark using the Equal Error Rate (EER) (Wu et al., 2015) metric on seven tasks using the LwF method. All methods and tasks are trained using only-1 linear layer model. The EER results are illustrated as follows:
EER Task1 Task2 Task3 Task4 Task5 Task6 Task7
Train after Task1 0.26 0.334 0.4027 0.428 0.428 0.4847 0.4857
Train after Task2 0.478 $\uparrow$ 0.4836 $\uparrow$ 0.4844 $\uparrow$ 0.4887 $\uparrow$ 0.4893 $\uparrow$ 0.4972 $\uparrow$ 0.522 $\uparrow$
Train after Task3 0.5539 $\uparrow$ 0.5818 $\uparrow$ 0.6064 $\uparrow$ 0.6282 $\uparrow$ 0.6478 $\uparrow$ 0.6654 $\uparrow$ 0.6813 $\uparrow$
Train after Task4 0.6504 $\uparrow$ 0.6221 $\uparrow$ 0.5962 $\downarrow$ 0.5723 $\downarrow$ 0.5503 $\downarrow$ 0.5299 $\downarrow$ 0.511 $\downarrow$
Train after Task5 0.4988 $\downarrow$ 0.496 $\downarrow$ 0.4972 $\downarrow$ 0.497 $\downarrow$ 0.4968 $\downarrow$ 0.4973 $\downarrow$ 0.4939 $\downarrow$
Train after Task6 0.495 $\downarrow$ 0.4949 $\downarrow$ 0.4949 $\downarrow$ 0.4956 $\downarrow$ 0.4959 $\downarrow$ 0.498 $\uparrow$ 0.497 $\uparrow$
Train after Task7 0.4922 $\downarrow$ 0.4915 $\downarrow$ 0.4914 $\downarrow$ 0.4915 $\downarrow$ 0.4909 $\downarrow$ 0.4972 $\downarrow$ 0.4984 $\uparrow$
Table 2: Training Performance After Each Task
Table 2 summarizes the training performance of the model after each task in a sequential learning scenario. The evaluation metric used is Equal Error Rate (EER), and tasks are denoted from Task1 to Task7. The table reveals the model’s performance in terms of EER after training on each task sequentially. Results show that initially, after Task1, the EER is 0.26, gradually improving to 0.4857 after Task7. Notably, training after Task2 leads to a significant improvement, marked with an upward arrow, and the trend continues until Task4. However, after Task4, there is a decline in performance (marked with downward arrows) which persists through Task7. This suggests that the model might start to overfit or lose generalization capabilities after learning from Task4 onwards. Additionally, Task6 shows a slight improvement, suggesting some recovery in performance. Overall, the table provides insights into how the model’s performance evolves over sequential tasks, indicating both improvements and potential challenges in continual learning scenarios.
(a) Task 1
(b) Task 2
(c) Task 3
(d) Task 4
(e) Task 5
(f) Task 6
(g) Task 7
Figure 1: Evaluation on Tasks

### v3

# Towards Robust Audio Deepfake Detection: A Evolving Benchmark for Continual Learning

@@ Abstract
The rise of advanced large language models such as GPT-4, GPT-4o, and the Claude family has made fake audio detection increasingly challenging. Traditional fine-tuning methods struggle to keep pace with the evolving landscape of synthetic speech, necessitating continual learning approaches that can adapt to new audio while retaining the ability to detect older types. Continual learning, which acts as an effective tool for detecting newly emerged deepfake audio while maintaining performance on older types, lacks a well-constructed and user-friendly evaluation framework. To address this gap, we introduce EVDA, a benchmark for evaluating continual learning methods in deepfake audio detection. EVDA includes classic datasets from the Anti-Spoofing Voice series, Chinese fake audio detection series. It supports various continual learning techniques, such as Elastic Weight Consolidation (EWC), Learning without Forgetting (LwF), and recent continual learning methods for audio deepfake detection like Regularized Adaptive Weight Modification (RAWM) and Radian Weight Modification (RWM). Additionally, EVDA facilitates the development of robust algorithms by providing an open interface for integrating new continual learning methods. Our code will be available at https://github.com/Cecile-hi/Evolving-FAD-CL-Benchmark.git.
@@ 1 Introduction
2 Related Work
2.1 Audio Deepfake Detection
The rise of synthetic audio, often referred to as deepfake audio, has led to significant advancements in detection techniques aimed at distinguishing between genuine and manipulated audio content. Early methods primarily focused on spectral analysis and traditional machine learning approaches, leveraging features such as Mel-frequency cepstral coefficients (MFCCs) and Gaussian Mixture Models (GMMs) to detect anomalies indicative of synthetic audio (Wu et al., 2015). With the evolution of deep learning, more sophisticated models, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), have been developed to capture subtle temporal and spectral differences between real and fake audio. Recent advancements have introduced hybrid models that combine deep learning with signal processing techniques to improve detection accuracy in challenging environments, such as noisy or low-quality recordings (Todisco et al., 2019; Liu et al., 2023). Additionally, cross-lingual and cross-dataset generalization has become a focal point, with new datasets and benchmarks being established to test models against a wider range of synthetic audio scenarios (Yi et al., 2022; Müller et al., 2022).
2.2 Continual Learning Methods
Continual learning, also known as lifelong learning, addresses the challenge of training models on sequential tasks while retaining knowledge from previous tasks. In the context of deepfake audio detection, this capability is crucial as new types of attacks and spoofing techniques emerge. Traditional models suffer from catastrophic forgetting, where learning new tasks leads to the erosion of previously acquired knowledge. To mitigate this, methods such as Elastic Weight Consolidation (EWC) have been proposed, which constrain the updates of important parameters to preserve knowledge from earlier tasks (Kirkpatrick et al., 2017). Other techniques, such as experience replay (Rolnick et al., 2019) and knowledge distillation (Li & Hoiem, 2017), have also been employed to maintain performance across tasks by rehearsing previous data or transferring knowledge from a stable teacher model to a student model during training. Recent advancements in continual learning have focused on adapting these methods to complex neural architectures, such as transformers and deep CNNs, making them more robust in dynamic environments where the nature of the data and tasks evolves over time (Rebuffi et al., 2017). In the context of deepfake detection, these continual learning approaches (Zhang et al., 2023; 2024) are pivotal in ensuring that models can adapt to new challenges without (or miner) losing the ability to detect previously known spoofing techniques.
@@ 3 Proposed Method
To address the growing complexity and diversity of audio deepfake detection, we propose a comprehensive benchmark designed to rigorously evaluate the performance of anti-spoofing models across a variety of challenging scenarios. Existing audio deepfake detection methods often focus on limited conditions or datasets, failing to capture the full range of real-world variations that models may encounter (Ma et al., 2021; Zhang et al., 2023; 2024). Our benchmark aims to fill this gap by incorporating eight distinct tasks, each carefully selected to represent different data sources, languages, and acoustic conditions. This diversity allows for a thorough assessment of model robustness, adaptability, and cross-lingual capabilities, pushing the boundaries of current audio spoofing detection methods. The benchmark is structured to challenge models with varying degrees of difficulty, ensuring that only the most advanced and resilient models can succeed across all tasks.
@@ 3.1 Data Architecture
Our benchmark architecture consists of eight tasks (Task 1 to Task 8), each representing different datasets and conditions for detecting synthetic audio. These tasks vary in data sources, language, acoustic conditions, and keep the number of training and evaluation samples in same. The data architecture of EVDA benchmark are listed as Table 1, which is designed to comprehensively evaluate the performance of audio spoofing and anti-spoofing models across a diverse set of conditions and datasets. Below, we detail each experiment’s setup and objectives:
Table 1: Data architecture of our proposed benchmark
ID Source Have Vocoder ID Language Acoustic Condition Train Test
Task 1 FMFCC-A F Chinese Human 2000 5000
Task 2 In-the-Wild F English Real world 2000 5000
Task 3 ADD 2022 F Chinese Low-quality/Partially fake 2000 5000
Task 4 ASVspoof2015 T English No significant noise 2000 5000
Task 5 ASVspoof2019LA T English Logical access 2000 5000
Task 6 ASVspoof2021LA T English Logical access 2000 5000
Task 7 FoR F English Human 2000 5000
Task 8 HAD F Chinese Human/Partially fake 2000 5000
@@
Task 1: FMFCC-A (Chinese)(Zhang et al., 2021) - Utilizes 2000 training and 5000 evaluation samples of Chinese audio. This experiment is critical for models to adapt to low-quality, partially fake audio environments, increasing the difficulty of the spoofing detection task.
@@
Task 2: In-the-Wild (English)(Müller et al., 2022) - Uses 2000 training and 5000 evaluation samples in real-world audio scenarios, testing the models’ effectiveness in dynamically changing and challenging acoustic environments.
Task 3: ADD 2022 (Chinese)(Yi et al., 2022) - Incorporates 2000 training and 5000 evaluation samples of Chinese audio. This experiment focuses on low-quality, partially fake settings, presenting challenges in detecting spoofed audio.
Task 4: ASVspoof2015 (English)(Wu et al., 2015) - Utilizes 2000 training and 5000 evaluation samples from the ASVspoof2015 dataset, focusing on English language audio in clean conditions without significant channel or background noise effects. Vocoder ID is available, aiding in more precise synthetic speech detection.
Task 5: ASVspoof2019 - Logical Access (LA) (English)(Todisco et al., 2019) - Incorporates 2000 training and 5000 evaluation samples focusing on logical access scenarios to test models against sophisticated spoofing attacks in English. Vocoder ID is again available, providing critical data for assessing synthetic signal traits.
Task 6: ASVspoof2021 - Logical Access (LA) (English)(Liu et al., 2023) - Continues the logical access examination with the latest dataset iteration, using English audio samples. The presence of vocoder ID helps in identifying new spoofing techniques that have emerged since the 2019 dataset.
Task 7: FoR (English)(Reimao & Tzerpos, 2019) - Uses 2000 training and 5000 evaluation samples of purely human-generated audio in English, testing the robustness of models against non-synthetic audio scenarios.
Task 8: HAD (Chinese)(Yi et al., 2021) - Utilizes 2000 training and 5000 evaluation samples of Chinese audio, which includes a mix of human and partially fake audio, presenting challenges in cross-lingual and authenticity discernment.
3.2 Arrangement of the tasks
We have arranged and highlighted the table data based on the distinct characteristics of each experiment on Table 1, emphasizing the differences in language and acoustic conditions. This organization allows for a clearer comparison of how various datasets and conditions affect the performance of deepfake audio detection models. By highlighting the changes, we can better understand the specific challenges each dataset introduces and how the models need to adapt to these variations. The key distinctions in the experiments are reflected in the changes to both language and acoustic conditions as follows:
- Task 1 to Task 2: The language changes from Chinese to English, with the acoustic condition shifting from ”Human” to ”Real world.” This transition highlights the model’s need to adapt from a controlled environment to more natural, varied conditions.
- Task 2 to Task 3: The language switches back to Chinese, while the acoustic condition changes to ”Low-quality/Partially fake,” indicating a mix of real and synthetic audio in more challenging quality conditions.
- Task 3 to Task 4: The language changes to English, with the acoustic condition becoming ”No significant noise,” focusing on clean audio without significant interference.
- Task 4 to Task 5: The language remains English, but the focus shifts to ”Logical access” scenarios, introducing sophisticated spoofing attacks.
- Task 5 to Task 6: The language and acoustic condition remain consistent (English, ”Logical access”), but the dataset iteration updates to the latest version with newly emerged spoofing technologies.
- Task 6 to Task 7: The language stays the same (English), while the acoustic condition returns to ”Human,” emphasizing purely human-generated audio.
- Task 7 to Task 8: The language changes back to Chinese, with the acoustic condition now being ”Human/Partially fake,” involving a mix of human and synthetic audio.
@@ 3.2 Arrangement of the tasks
These highlighted changes are crucial for understanding the adaptability and robustness of deepfake audio detection models across different datasets, languages, and acoustic conditions. The benchmark is designed not only to assess the current state-of-the-art in spoofing detection but also to push the development of more robust, adaptable, and cross-lingual models capable of handling the complexities of real-world audio spoofing scenarios.
@@ 3.3 Data Selection Methodology
For each task within our benchmark, we provided two strategies for data selection:
Random Selection: We randomly select a fixed number of authentic and spoofed audio samples from the original datasets or collected data. Initially, we split the data into two groups: Real and Fake. From the ”Real” group, we randomly select 1000 real samples, and from the ”Fake” group, we select 1000 fake samples corresponding to various generation methods. These samples are combined to form the training set. For the evaluation set, we follow the same procedure but select only 2500 real and 2500 fake samples. This approach ensures a balanced representation of real and fake samples in both the training and evaluation sets. In current benchmark, this method has been used for data selection.
@@
Here, $N_{\text{real}}$ and $N_{\text{fake}}$ represent the number of models classifying the sample as real or fake, respectively. According to information theory (Shannon, 1948), samples with the highest entropy values contain the most information, indicating uncertainty and hence, difficulty in classification. We list the entropy of each sample and rank them from highest to lowest. The top 2500 real and 2500 fake samples are selected to construct the evaluation set, ensuring it includes the most challenging and informative samples. The next 1000 real and 1000 fake samples are selected for the training set, which benefits from the inclusion of diverse and informative samples.
3.4 Continuous Learning Methods in proposed Benchmark
The benchmark evaluates various continual learning methods to enhance the adaptability of audio spoofing detection systems. These approaches are summarized as follows:
Replay (Rolnick et al., 2019) mitigates catastrophic forgetting by replaying a subset of previously seen samples during training, helping to maintain performance on earlier tasks.
Finetuning adapts a pre-trained model to new tasks by continuing training with new data, effectively leveraging existing models to incorporate new knowledge.
EWC (Kirkpatrick et al., 2017) adds a regularization penalty to the loss function based on parameter importance, preserving critical knowledge from previous tasks.
GDumb (Prabhu et al., 2020) combines finetuning with periodic retraining on a data subset, balancing new and old information for large-scale learning.
CWRStar (Maltoni & Lomonaco, 2019) employs rehearsal and interleaving tasks using a fixed memory, improving retention of past knowledge while learning new tasks.
SI (Zenke et al., 2017) introduces dynamic regularization based on the importance of model parameters, balancing the retention of old knowledge with the acquisition of new information.
OWM (Zeng et al., 2019) adjusts weight updates to remain orthogonal to those associated with previous tasks, reducing forgetting by preserving relevant features.
RAWM (Zhang et al., 2023) adapts weight updates based on the ratio of genuine to fake utterances, introducing regularization to retain knowledge across different acoustic conditions.
RWM (Zhang et al., 2024) uses radian-based weight modification to minimize disruption to previously learned features, enhancing stability and adaptability in continual learning.
In summary, our benchmark provides a comprehensive evaluation of these methods, aiming to develop robust models that can effectively handle evolving audio spoofing attacks with high reliability and accuracy.
@@ 4 Result
4.1 Compared with multiply continual learning methods
Methods Task1 Task2 Task3 Task4 Task5 Task6 Task7 Task8 Avg
Replay 2.88 7.60 2.56 1.40 2.80 8.84 6.76 4.12 4.62
Finetuning 8.28 10.04 5.36 1.52 4.08 18.24 10.44 4.44 7.80
EWC 6.64 10.16 5.52 1.72 4.20 16.92 10.08 4.56 7.475
GDumbFinetune 2.88 8.28 4.92 3.64 5.88 19.28 13.16 10.44 8.56
CWRStar 28.2 36.72 22.48 29.8 46.4 49.36 47.04 34.12 36.77
SI 8.52 10.2 7.48 1.68 5.40 17.12 13.92 4.60 8.615
OWM 6.72 28.52 32.88 14.32 19.84 40.8 26.08 21.88 23.88
RAWM 9.24 13.6 5.38 1.68 4.36 16.4 12.44 6.00 8.638
RWM 9.64 14.76 8.52 2.04 4.64 17.6 11.88 5.96 9.38
Table 2: EER (%) on each task for different continual learning methods. The cells with the lowest EER in each column are highlighted in red, while the cells with the second-lowest EER are highlighted in yellow. Lower EER values indicate better performance.
Table 2 summarizes the performance of various continual learning methods across eight tasks, evaluated by Equal Error Rate (EER) (Wu et al., 2015). Compared with all continual learning methods, replay consistently shows most competitive performance across all tasks, demonstrating its overall effectiveness. Finetuning and EWC exhibit relatively similar performances. However, EWC achieves the better performance in old and average tasks, indicating its overcoming forgetting performance in old tasks. GDumbFinetune performs well in certain tasks, notably achieving the lowest EER in Task 1 along with Replay. However, its overall performance is less consistent, with a higher average EER. CWRStar consistently shows higher EER values across all tasks, indicating it is less effective compared to other methods, with the highest average EER (36.77).SI and OWM demonstrate mixed results. SI achieves low EERs in some tasks, but not as consistently as Replay or EWC. OWM shows the highest EER in several tasks, particularly in Task2 and Task3. RAWM and RWM provide relatively competitive performances but fall behind Replay and EWC in several tasks. RAWM has a notable second-lowest EER in Task 6 compared to the best. Overall, Replay and EWC are the top-performing methods, with Replay showing the lowest average EER. These results suggest that Replay and EWC are effective in minimizing error rates in continual learning scenarios for audio deepfake detection, while other methods like CWRStar and OWM may need further refinement for better performance.
4.2 Continual learning on each task
@@
We evaluate our benchmark using the EER metric on 8 tasks using the conventional Finetuning and EWC continual learning method. All methods and tasks are trained using 5 linear layers model with 128-hidden dimension. The EER results are illustrated as Table 3:
Train after Task1 Task2 Task3 Task4 Task5 Task6 Task7 Task8
Task1 FT 5.44 32.12 28.52 33.80 42.12 39.20 45.00 22.32
EWC 5.44 32.12 28.52 33.80 42.12 39.20 45.00 22.32
Task2 FT 10.20 6.32 25.84 7.72 10.64 30.60 21.92 12.28
EWC 9.48 6.20 26.72 8.96 12.24 29.92 23.00 12.24
Task3 FT 11.48 15.04 1.48 19.16 44.40 34.92 46.28 19.72
EWC 12.92 15.36 1.64 20.40 49.80 36.84 42.76 20.16
Task4 FT 16.56 13.00 4.12 1.80 6.32 25.76 17.36 9.92
EWC 17.96 12.80 4.80 1.72 6.16 27.32 16.76 9.92
Task5 FT 18.56 10.28 9.28 1.76 2.64 23.60 12.08 11.08
EWC 21.04 9.96 10.60 1.68 2.36 23.88 12.16 10.96
Task6 FT 10.04 12.36 5.64 2.40 6.96 9.96 14.12 11.16
EWC 10.12 11.52 6.20 2.16 6.52 10.04 13.84 11.52
Task7 FT 10.84 12.72 8.20 1.72 2.64 12.44 6.64 12.40
EWC 11.08 11.96 9.72 2.04 2.52 12.84 6.28 12.64
Task8 FT 8.28 10.04 5.36 1.52 4.08 18.24 10.44 4.44
EWC 6.64 10.16 5.52 1.72 4.20 16.92 10.08 4.56
Table 3: The EER(%) evaluated on all tasks after each training using conventional finetuning (FT) and EWC.
The results presented in Table 3 compare the EER (%) across all tasks after each training session using two methods: conventional finetuning (FT) and Elastic Weight Consolidation (EWC). From the table, it is evident that both methods exhibit similar performance when applied to the first task, as expected, since no prior knowledge needs to be preserved. However, as training progresses on subsequent tasks, noticeable differences emerge between the two methods. For Task 2, EWC shows slightly improved performance over FT in both Task 1 and Task 2, with reductions in EER ranging from minor differences. These improvements suggest that EWC effectively mitigates catastrophic forgetting by preserving knowledge from earlier tasks, allowing for more stable learning. However, there are cases, like Task 3, where EWC does not offer significant improvements over FT, indicating that its effectiveness can be task-dependent. Overall, while both methods show varying levels of performance across different tasks, EWC generally provides a modest improvement over FT on old tasks. This highlights advantage in continual learning where the model needs to maintain performance across multiple tasks without revisiting them.
5 Discussion
In this work, we have proposed a continual learning benchmark, which is the first as best as we know, for audio deepfake detection. One of the key advantages of our work is the thorough evaluation of continual learning methods in a challenging domain-shift condition. The varied and complex nature of audio data, combined with the evolving techniques used in generating deepfakes, makes it a particularly demanding area for continual learning. Our results provide valuable insights into the current capabilities of these methods and their applicability to real-world problems.
However, our study also has certain limitations. Firstly, the scope of continual learning methods explored in this work is limited. Although we have focused on well-established methods like FT and EWC, there are several other approaches that could offer additional benefits, such as replay-based methods, regularization techniques, and meta-learnin approaches. Furthermore, the benchmark tasks utilized in this study, while representative, may not fully capture the diversity of audio deepfakes encountered in practice.
In future work, we plan to expand our investigation to include a broader range of continual learning methods. This will not only provide a more complete understanding of the strengths and weaknesses of different approaches but also enable the proposal of novel methods tailored specifically to the challenges of audio deepfake detection. Additionally, we intend to incorporate newly emerging audio deepfake techniques, such as those powered by advanced generative models like GPT-4 and its successors, into our evaluation framework. By doing so, we aim to ensure that our models remain robust and effective in detecting the latest deepfake techniques, thereby maintaining their relevance in the face of rapidly advancing technologies.